OpenStudy (anonymous):

what is the name of the polar graph r=2? and a characteristic of it (radius, location, point on the graph, etc.)?

myininaya (myininaya):

its a circle! :)

OpenStudy (anonymous):

radius is two. you get a circle.

jimthompson5910 (jim_thompson5910):

This is a circle of radius 2. Basically the radius is fixed because r = 2

jimthompson5910 (jim_thompson5910):

Oh and it's centered at (0,0)

jimthompson5910 (jim_thompson5910):

should be x^2+y^2=r^2, so x^2+y^2=2^2 or x^2+y^2=4